# Freeze-in Gravitational Waves and Dark Matter in Warm Inflation

### v3

@@ Abstract
Recent study [1] has suggested that warm inflation may be realized with Standard Model gauge interaction alone. Motivated by this framework, we investigate the gravitational wave spectrum and graviton-portal dark matter production through the freeze-in process generated during warm inflation scenarios. We perform a comparative analysis for different dissipation terms, focusing on their distinct gravitational wave signatures in the high-frequency regime. Our findings reveal qualitative and quantitative differences in the spectral behavior, offering a preliminary pathway for exploring inflationary and dark matter models through high-frequency gravitational wave signals.
@@ 1 Introduction
Cosmic inflation [2, 3, 4] can naturally explain the cosmic microwave background radiation and the large-scale structure observations, since the density perturbation of the inflation field can provide the primordial seeds of our Universe. Warm inflation (WI) [5] is one type of inflationary scenario that fundamentally alters the conventional picture of the early universe’s expansion by including dissipative effects and concurrent radiation production during inflation. Unlike the standard cold inflation (CI) model, where the universe supercools and reheating occurs from the late stage of inflation [6], WI maintains a thermal bath throughout the inflation process by dissipating the inflaton field’s energy into radiation. This continuous particle production modifies the dynamics of the inflation field and the generation of primordial density perturbations, offering possible solutions to challenges faced by cold inflation such as the “eta problem” and allowing for sub-Planckian field excursions. The conceptual foundation of WI integrates quantum field theory techniques to model dissipative processes, showing that interactions between the inflation field and other fields produce radiation simultaneously with inflationary expansion.
@@
Recently, K.V. Berghaus, M. Drewes, and S. Zell in ref. [1],demonstrate, for the first time, that WI can be consistently realized through Standard Model gauge interactions, and the CMB predictions of this model have been further investigated in [8]. This finding provides an intriguing framework for dissipative dynamics during inflation and establishes a concrete connection between early-universe cosmology and well-motivated particle physics interactions. Motivated by this idea in ref. [1],we investigate the production of gravitational wave (GW) originating from the scattering of thermal plasma constituents and the graviton-portal dark matter (DM) during the inflationary epoch and estimate their relic density. The GW produced during reheating in the standard cold inflation paradigm has been investigated before [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In WI, the universe naturally transitions from the inflationary epoch to the radiation-dominated era without requiring a separate reheating phase, and gravitons can be continuously produced from plasma [22, 23, 24, 25, 26, 27]. The graviton and DM production mechanism closely resembles the freeze-in DM scenarios [28, 29, 30]. In this work, we mainly investigate the GW spectrum in the high-frequency regime and the associated DM production mediated by gravitons, within the framework of WI. Our analysis encompasses both a simplified version of the recently proposed WI scenario based solely on Standard Model gauge interactions [1], as well as the typical WI model considered in ref. [27].
This paper is organized as follows. In section 2, we briefly review the mechanism of WI. In section 3, we analyze the freeze-in mechanism of gravitons and DM in WI. In section 4, we calculate the GW spectrum for different WI models. In section 5, we also estimate DM production solely by gravitational interaction from Standard Model plasma, and study the correlation between the DM mass and the peak amplitude of GW spectra. Finally, a concise conclusion is given in section 6.
@@ 2 Warm inflation
During the WI, the evolution of the background is governed by the following equations:
@@
where $\phi$ is the inflaton field, $H$ the Hubble parameter, $\rho_{r}$ the radiation energy density, and $M_{\rm pl}\simeq 1.221\times 10^{19}\penalty 10000\ \mathrm{GeV}$ the Planck mass, respectively. $\Upsilon$ is the dissipation coefficient which decides the energy transfer between the inflaton and the thermal bath. The distinction between the warm and cold inflation is that in the former model, the dissipation term causes the radiation energy not to vanish. In this work, we assume the inflation potential
@@
For convenience, one defines the dimensionless dissipation ratio $Q\equiv\frac{\Upsilon}{3H}$, which gives a measure of the strength of the dissipation effects in comparison to the Hubble expansion. Then, we can classify WI based on the parameter $Q$ as follows: for $Q>1$, strong dissipative regime; for $Q<1$, weak dissipative regime.
@@
The dissipation coefficient $\Upsilon$ has the general form
$\Upsilon\approx\kappa\frac{T^{a}\phi^{b}}{f^{a+b-1}}$ (2.8)
where $\kappa$ is a constant. $a$, and $b$ could be some numerical powers. $f$ could be the symmetry-breaking scale. There are some models that provide the specific form of $\Upsilon$. For example, refs. [31, 32, 1] realize the WI by introducing axion coupling to a pure Yang-Mills gauge group and give $\Upsilon\propto T^{3}$. Ref. [33] proposes a scenario which is similar to the little Higgs model, where $\Upsilon\propto T$.
In this work, we discuss these two cases separately, $\Upsilon\propto T$, $T^{3}$, and compare their corresponding predictions. For each case, we consider $Q_{\rm ini}=0.01$ and $Q_{\rm ini}=1$, where $Q_{\rm ini}$ denotes the value of $Q$ at the beginning of inflation. Using the WI2easy code [34], we have verified that both values satisfy the CMB constraints from Planck data [35].
3 Freeze-in production in warm inflation: gravitons and dark matter
In WI, gravitons are continuously produced from the plasma via particle scattering processes during and after inflation. Since gravitons are produced out of equilibrium and subject to Planck-scale suppression, their production mechanism is conceptually very similar to the ultra-violet DM freeze-in production mechanism [30]. And in ref. [28], the authors found that freeze-in-produced DM relic density could be significantly enhanced in WI compared with CI. In this section, we argue that this enhancing mechanism applies equally to freeze-in production of gravitons.
In this work, we only consider gravitons with sufficiently high momenta to ensure that they always remain within the horizon. Their evolution is governed by the Boltzmann equation
@@ 3 Freeze-in production in warm inflation: gravitons and dark matter
where $\rho_{\rm GW}$ is the GW energy density. $\gamma$ corresponds to the production rate of graviton energy density from the SM thermal bath and it is a function of temperature $T$. Using dimensional analysis, we obtain
@@
where $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ are coefficients. The first term corresponds to the production of a single graviton. Since $T\ll M_{\rm pl}$ during WI as we will show later, we consider only the leading-order contribution. For double-graviton production, see refs. [30, 26].
@@
Using $dN_{e}/dt=H$, the left-handed side of eq. (3.1) can be re-parametrized as
$\frac{d\rho_{\mathrm{GW}}}{dt}+4H\rho_{\mathrm{GW}}=He^{-4N_{e}}\frac{d(\rho_{\mathrm{GW}}e^{4N_{e}})}{dN_{e}}$ (3.3)
Then the solution to eq. (3.1) reads
$\rho_{\rm GW}=e^{-4N_{e}}\int_{N_{e}^{i}}^{N_{e}}e^{4N_{e}^{\prime}}\frac{\mathcal{\gamma}}{H}\ dN_{e}^{\prime}$ (3.4)
where we have assumed that the energy density of GW is zero at the onset of inflation.
Now, the production rate of the comoving GW energy density, $I_{\mathrm{GW}}=d(e^{4N_{e}}\rho_{\mathrm{GW}})/dN_{e}$, reads
$I_{\mathrm{GW}}(N_{e})=e^{4N_{e}}\frac{\mathcal{\gamma}}{H}=\mathcal{C}_{1}e^{4N_{e}}\frac{T^{7}(N_{e})}{M_{\rm pl}^{2}H(N_{e})}$ (3.5)
@@
From eq. (3.5), during the inflation stage, we can see that $I_{\mathrm{GW}}$ is an exponentially increasing function. This behavior arises because the inflaton energy dominates the universe, as well as the temperature $T$ and the Hubble parameter $H$ remain approximately constant during the inflation stage.
And when the universe transits to the radiation domination stage, $I_{\mathrm{GW}}$ will be proportional to $e^{-N_{e}}$ since $T\propto e^{-N_{e}}$ and $H\propto e^{-2N_{e}}$. Therefore, $I_{\mathrm{GW}}$ should sharply peaked at some e-fold, $N_{\mathrm{peak}}^{\mathrm{GW}}$, which satisfies
$\left.\left[4+7\frac{d\ln{T(N_{e})}}{dN}-\frac{d\ln{H(N_{e})}}{dN}\right]\right|_{N_{e}=N_{\mathrm{peak}}^{\mathrm{GW}}}=0$ (3.6)
This is similar to the case of DM freeze-in process, where the DM yield reads [28]
$Y_{\mathrm{DM}}(N_{e})=\frac{45}{2\pi^{2}g_{*,s}}\frac{e^{-3N_{e}}}{T^{3}(N_{e})}\int_{N_{e}^{i}}^{N_{e}}I_{\mathrm{DM}}(N_{e}^{\prime})\ dN_{e}^{\prime}$ (3.7)
And $I_{\mathrm{DM}}$ is the production rate of the comoving DM number density
$I_{\mathrm{DM}}(N_{e})=e^{3N_{e}}\frac{T^{2n+4}(N_{e})}{\Lambda^{2n}H(N_{e})}$ (3.8)
where $\Lambda$ is the cutoff of the effective field theory which describes the interaction between DM and the radiation. For similar reason, the DM production should be sharply peaked at $N_{\mathrm{peak}}^{\mathrm{DM}}$, and it satisfies
$\left.\left[3+(2n+4)\frac{d\ln{T(N_{e})}}{dN_{e}}-\frac{d\ln H(N_{e})}{dN_{e}}\right]\right|_{N_{e}=N_{\mathrm{peak}}^{DM}}=0$ (3.9)
Figure 1: The evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI with dissipation term $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, respectively. All curves are normalized to their respective maximum values. The blue line corresponds to $Q_{\mathrm{ini}}=0.01$ and the red line corresponds to $Q_{\mathrm{ini}}=1$. The black dashed line indicates the end of inflation ($\epsilon_{H}=1$), while the red and blue dashed lines correspond to the onset of radiation domination ($\epsilon_{H}=2$), respectively. Note that $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ both sharply peaked at some e-folding numbers in the post-inflationary epoch, prior to the onset of radiation domination. It can be also found that the reheating process is much more efficient for $\Upsilon\propto T^{3}$ than for $\Upsilon\propto T$.
Figure 1 shows the evolution of $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ in WI for both $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, under both $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$. We set $n=2$ since this is the case of purely gravitational DM as we will discuss later. All curves are normalized to their respective peak values. The blue and red solid lines represent the cases with $Q_{\mathrm{ini}}=0.01$ and $Q_{\mathrm{ini}}=1$, respectively. The black dashed line marks the end of inflation, while the blue and red dashed lines indicate the onset of radiation domination for each scenario. Notably, both $I_{\mathrm{DM}}$ and $I_{\mathrm{GW}}$ exhibit obvious peaks at specific e-folding numbers during the post-inflationary epoch, preceding the transition to radiation domination.
Now, the DM yield can be expressed approximately as [28]
$Y_{\rm DM}(N_{e})\simeq\frac{45}{2\pi^{2}g_{*s}}\frac{e^{3(N_{\mathrm{peak}}^{\rm DM}-N_{e})}}{\Lambda^{2n}T^{3}(N_{e})}\Delta N_{\mathrm{peak}}^{\rm DM}\times\frac{T^{2n+4}(N_{\mathrm{peak}}^{\rm DM})}{H(N_{\mathrm{peak}}^{\rm DM})}\quad(N_{e}>N_{\mathrm{peak}}^{\rm DM})$ (3.10)
@@
with $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$ being the half-width of $I_{\rm DM}(N_{e})$.
In CI, DM can be produced during the radiation-dominant epoch which starts at some initial time $N_{e}^{\rm RH}$ with the reheating temperature $T_{\mathrm{RH}}$. The DM yield then reads
$Y_{\rm DM}^{\mathrm{RH}}\simeq\frac{1}{4\sqrt{\pi}}\left(\frac{45}{\pi^{2}g_{*}}\right)^{\frac{3}{2}}\frac{1}{2n-1}\frac{M_{\rm pl}T_{\mathrm{RH}}^{2n-1}}{\Lambda^{2n}}$ (3.11)
In WI, the reheating temperature is defined as the plasma temperature when the universe becomes radiation-dominated ($\epsilon_{H}=2$). For identical reheating temperatures, we can compare the DM relic abundance from WI and CI,
$\frac{Y_{\rm DM}(N_{e})}{Y_{\rm DM}^{\mathrm{RH}}}\simeq(2n-1)\frac{I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})}{I_{\rm DM}(N_{e}^{\mathrm{RH}})}\Delta N_{\mathrm{peak}}^{\rm DM}$ (3.12)
Since $I_{\rm DM}(N_{\mathrm{peak}}^{\rm DM})\gg I_{\rm DM}(N_{e}^{\rm RH})$ and $\Delta N_{\mathrm{peak}}^{\rm DM}\gtrsim 1$, the relic abundance is significantly enhanced in WI compared with CI. Same argument should apply to the freeze-in GW. Analogously to freeze-in DM, it is useful to define the dimensionless quantity $\rho_{\rm GW}/s^{\frac{4}{3}}$ with $s=\frac{2\pi^{2}g_{*s}}{45}T^{3}$ the entropy density [22]
@@
$\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{e^{4(N_{\mathrm{peak}}^{\rm GW}-N_{e})}}{M_{\rm pl}^{2}T^{4}(N_{e})}\Delta N_{\mathrm{peak}}^{\rm GW}\times\frac{T^{7}(N_{\mathrm{peak}}^{\rm GW})}{H(N_{\mathrm{peak}}^{\rm GW})}\quad(N_{e}>N_{\mathrm{peak}}^{\rm GW})$ (3.13)
where $\Delta N_{\mathrm{peak}}^{\rm GW}\gtrsim 1$ is the half-width of $I_{\rm GW}(N_{e})$.
@@
For CI with the same reheating temperature $T_{\mathrm{RH}}$, by integrating eq. (3.4) from $N_{e}^{\mathrm{RH}}$ to infinity we can obtain
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI},T_{\rm RH}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{T_{\rm RH}^{3}}{M_{\rm pl}^{2}H_{\rm RH}}$ (3.14)
where $H_{\rm RH}$ is the Hubble parameter at $N_{e}^{\rm RH}$.
@@
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]\bigg/\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI},T_{\rm RH}}\simeq\frac{I_{\rm GW}(N_{\rm peak}^{\rm GW})}{I_{\rm GW}(N_{e}^{\rm RH})}\Delta N_{\mathrm{peak}}^{\rm GW}$ (3.15)
We conclude that GW are primarily generated around $N_{\rm peak}^{\rm GW}$, with their production enhanced in WI. Moreover, we expect a more pronounced enhancement for the WI model with $\Upsilon\propto T$ than for the $\Upsilon\propto T^{3}$ model, as the more efficient reheating process in the latter case leads to a significantly smaller value of the ratio $I_{\rm GW}(N^{\rm GW}_{\rm peak})\big/I_{\rm GW}(N^{\rm RH}_{e})$.

### v4

@@ Abstract
Recent study [1] has suggested that warm inflation may be realized with a minimal extension of the Standard Model by a single scalar inflaton field with an axion-like coupling to gluons. Motivated by this framework, we investigate the gravitational wave spectrum and graviton-portal dark matter production through the freeze-in process generated during warm inflation scenarios. We perform a comparative analysis for different dissipation terms, focusing on their distinct gravitational wave signatures in the high-frequency regime. Our findings reveal qualitative and quantitative differences in the spectral behavior, offering a preliminary pathway for exploring inflationary and dark matter models through high-frequency gravitational wave signals.
@@ 1 Introduction
Cosmic inflation [2, 3, 4] can naturally explain the cosmic microwave background radiation and the large-scale structure observations, since the density perturbation of the inflaton field can provide the primordial seeds of our Universe. Warm inflation (WI) [5, 6, 7, 8, 9] is one type of inflationary scenario that fundamentally alters the conventional picture of the early Universe’s expansion by including dissipative effects and concurrent radiation production during inflation. Unlike the standard cold inflation (CI) model, where the Universe supercools and reheating occurs from the late stage of inflation [10], WI maintains a thermal bath throughout the inflation process by dissipating the inflaton field’s energy into radiation. This continuous particle production modifies the dynamics of the inflaton field and the generation of primordial density perturbations, offering possible solutions to challenges faced by CI such as the “eta problem” and allowing for sub-Planckian field excursions. The conceptual foundation of WI integrates quantum field theory techniques to model dissipative processes, showing that interactions between the inflaton field and other fields produce radiation simultaneously with inflationary expansion.
@@
Recently, K.V. Berghaus, M. Drewes, and S. Zell in ref. [1], demonstrate, for the first time, that WI might be consistently realized within a minimal extension of the Standard Model, incorporating a scalar inflaton field with an axion-like coupling to gluons and a monomial potential. The CMB predictions and viability of this model have been further investigated in [12]. This finding might provide a possible framework for dissipative dynamics during inflation and establishes a concrete connection between early-Universe cosmology and well-motivated particle physics interactions. Motivated by this idea in ref. [1], we investigate two cosmic relics produced in WI scenario: (i) the gravitational waves (GWs) originating from the scattering of thermal plasma constituents, and (ii) the graviton-portal dark matter (DM). Among the many potential sources of primordial GWs [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], GWs generated by scattering processes among thermal plasma particles [36, 37, 38, 39, 40, 41, 42] are guaranteed to be present since it relies solely on the presence of a thermal bath. In this work, we refer to this GWs as “freeze-in graviton” or “freeze-in GWs” [39] since the graviton and the graviton-portal DM production mechanism closely resembles the freeze-in DM scenarios [43, 44, 45, 39]. The production of freeze-in particles is primarily determined by the thermal history of the early Universe; we therefore anticipate that it can effectively probe WI models, whose thermal history differs significantly from that of CI. In WI, the Universe naturally transitions from the inflationary epoch to the radiation-dominated era without requiring a separate reheating phase, allowing for the continuous production of gravitons and DM. In this work, we mainly investigate the freeze-in GW spectrum in the high-frequency regime and the associated DM production mediated by gravitons, within the framework of WI. Our analysis encompasses both a simplified version of the recently proposed WI scenario based solely on Standard Model gauge interactions [1], as well as the typical WI model considered in ref. [42].
This paper is organized as follows. In section 2, we briefly review the mechanism of WI. In section 3, we analyze the freeze-in mechanism of gravitons in WI, with a review of the analogous DM WI freeze-in provided in appendix A. In section 4, we calculate the GW spectrum for different WI models. In section 5, we also estimate DM production solely by gravitational interaction from Standard Model plasma, and study the correlation between the DM mass and the peak amplitude of GW spectra. Finally, a concise conclusion is given in section 6.
@@ 2 Warm inflation
In this section, we briefly review the background dynamics for general WI models. The evolution of the background is governed by the following equations:
@@
where $\phi$ is the inflaton field, $H$ the Hubble parameter, $\rho_{r}$ the radiation energy density, and $M_{\rm pl}\simeq 1.221\times 10^{19}\penalty 10000\ \mathrm{GeV}$ the Planck mass, respectively. The dot notation refers to the derivative with respect to cosmic time $t$, while the prime refers to the derivative with respect to the inflaton field $\phi$. Here, $\Upsilon$ is the dissipation coefficient which decides the energy transfer between the inflaton and the thermal bath. The distinction between the WI and CI is that in the former model, the dissipation term causes the radiation energy not to vanish. In this work, we assume the inflation potential
@@
For convenience, one defines the dimensionless dissipation ratio $Q\equiv\frac{\Upsilon}{3H}$, which gives a measure of the strength of the dissipation effects in comparison to the Hubble expansion. Then, we can classify WI based on the parameter $Q$ as follows: $Q>1$ marks the strong dissipative regime while $Q<1$ marks the weak dissipative regime.
@@
The specific form of the dissipation coefficient $\Upsilon$ depends on the microscopic physics underlying the WI model. This coefficient is generally a function of infalton $\phi$ and the radiation temperature $T$: $\Upsilon\equiv\Upsilon(\phi,T)$. For many WI models, the derived $\Upsilon(\phi,T)$ can be parameterized in the following form [9]:
$\Upsilon=C_{\Upsilon}T^{c}\phi^{p}M^{1-p-c}$ (2.8)
where $C_{\Upsilon}$ is a dimensionless constant, with $c$ and $p$ being numerical powers and $M$ a mass scale in the model. In this work, we focus on two representative and well-motivated WI models, which give the specific forms of the dissipation coefficients.
(i). The axion-like WI model introduces an axion-like coupling of the inflaton to a pure Yang-Mills gauge group, described by the following Lagrangian
$\mathcal{L}_{\rm int}=\frac{\alpha_{g}}{8\pi}\frac{\phi}{f_{a}}\tilde{F}^{a\mu\nu}F_{\mu\nu}^{a}$ (2.9)
where $F^{a}_{\mu\nu}$ is the field strength tensor and $\tilde{F}^{a\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F^{a}_{\rho\sigma}$. $\alpha_{g}\equiv g^{2}/4\pi$ with $g$ being the gauge coupling and $f_{a}$ corresponds to the axion decay constant in axion-like models. This interaction triggers sphaleron processes in the thermal bath, leading to a dissipation coefficient of the form [46, 47, 1].
$\Upsilon(T)=\frac{\Gamma_{\rm sph}(T)}{2f_{a}^{2}T}$ (2.10)
where $\Gamma_{\rm sph}(T)$ is the sphaleron rate which scales as $\Gamma_{\rm sph}(T)\sim\alpha_{g}^{5}T^{4}$ [48]. So one can obtain $\Upsilon\propto T^{3}$.
(ii). The warm little inflation model [49] proposes a scenario which is similar to the little Higgs model. The inflaton is a pseudo-Nambu-Goldstone boson of a broken gauge symmetry. The dissipation coefficient is derived as $\Upsilon\propto T$ from the inflaton’s coupling to light fermions.
We will discuss these two cases separately, $\Upsilon\propto T$, $T^{3}$, and compare their corresponding predictions. For each case, we consider $Q_{\rm ini}=0.01$ and $Q_{\rm ini}=1$, where $Q_{\rm ini}$ denotes the initial value of $Q$ at the beginning of inflation. Using the WI2easy code [50], we have verified that both values satisfy the CMB constraints from Planck data [51].
3 Freeze-in production of gravitons in warm inflation
In WI, gravitons are continuously produced from the plasma via particle scattering processes during and after inflation. Since gravitons are produced out of equilibrium and subject to Planck-scale suppression, their production mechanism is conceptually very similar to the ultraviolet (UV) DM freeze-in production mechanism [39]. And in ref. [44], the authors found that freeze-in-produced DM relic density could be significantly enhanced in WI compared with CI (a brief review is provided in appendix A). This conceptual similarity suggests that the freeze-in production of gravitons should be similarly enhanced in WI, which we focus on in this section.
In this work, we only consider gravitons with sufficiently high momenta to ensure that they always remain within the horizon. Their evolution is governed by the Boltzmann equation [22]
$\frac{\partial f_{h}}{\partial t}-kH\frac{\partial f_{h}}{\partial k}=C_{h}(k,T)$ (3.1)
where $f_{h}=f_{h}(t,k)$ and $k$ represent the distribution function of gravitons and graviton momentum, respectively. $C_{h}$ is the collision term for graviton production. For freeze-in process, the collision term is a function of the plasma temperature and graviton momentum, but independent of $f_{h}$ since the gravitons can never reach equilibrium and the backreactions from gravitons can be safely neglected.
The GW energy density is given by
$\rho_{\rm GW}=g_{h}\int\frac{d^{3}\bm{k}}{(2\pi)^{3}}kf_{h}$ (3.2)
with $g_{h}=2$ the degrees of freedom of the graviton. Now, by integrating eq. (3.1), we get evolution equation for GW energy density:
@@ 3 Freeze-in production of gravitons in warm inflation
where
$\gamma\equiv g_{h}\int\frac{d^{3}\bm{k}}{(2\pi)^{3}}kC_{h}(k,T)=g_{h}\int\frac{dk}{2\pi^{2}}k^{3}C_{h}(k,T)$ (3.4)
Now $\gamma$ corresponds to the graviton production rate density from the SM thermal bath and is a function of temperature $T$. Note that $\gamma$ has mass dimension 5. The leading-order contribution to $\gamma$ arises from single-graviton production processes (for example, the scattering process $ab\rightarrow ch$ where $a,b,c$ are SM particles and $h$ the graviton). The scattering amplitude for these processes is $\propto 1/M_{\rm pl}$, thereby suppressing the collision term—and hence $\gamma$ itself at leading order—by $1/M^{2}_{\rm pl}$. It then follows from dimensional analysis that the leading order term must scale as $T^{7}/M_{\rm pl}^{2}$. The next-to-leading order term, originating from double-graviton production [41], is further suppressed by $1/M_{\rm pl}^{4}$, and thus scales as $T^{9}/M_{\rm pl}^{4}$. Therefore, via dimensional analysis, we obtain:
@@
{where $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ are coefficients which depend on the couplings of the thermal bath particles appearing in the collision term. For simplicity, we ignore the running of the couplings and assume they are constants. A detailed calculation of the collision term can be found in refs. [37, 39, 41]. Here, we employ only the results of the dimensional analysis given in eq. (3.5), which suffices to provide preliminary insights into graviton production in WI. Moreover, since $T\ll M_{\rm pl}$ during WI as will be shown later, we consider only the leading-order contribution.
@@
Using $dN\equiv Hdt$, where $N$ is the number of e-folds, the left-handed side of eq. (3.3) can be re-parametrized as
$\frac{d\rho_{\mathrm{GW}}}{dt}+4H\rho_{\mathrm{GW}}=He^{-4N}\frac{d(\rho_{\mathrm{GW}}e^{4N})}{dN}$ (3.6)
Then the solution of eq. (3.6) reads
$\rho_{\rm GW}=e^{-4N}\int_{N_{0}}^{N}I_{\rm GW}(N^{\prime})\ dN^{\prime}$ (3.7)
where
$I_{\rm GW}(N)\equiv e^{4N}\frac{\mathcal{\gamma}}{H}=\mathcal{C}_{1}e^{4N}\frac{T^{7}(N)}{M_{\rm pl}^{2}H(N)}$ (3.8)
Note that $I_{\rm GW}$ is just the production rate of the comoving GW energy density, $e^{4N}\rho_{\rm GW}$, i.e. $I_{\mathrm{GW}}=d(e^{4N}\rho_{\mathrm{GW}})/dN$. And in deriving eq. (3.7), we have assumed a vanishing GW energy density at the initial time.
@@
From eq. (3.8), during the inflation stage, we can see that $I_{\mathrm{GW}}$ is an exponentially increasing function. This behavior arises because the inflaton energy dominates the Universe, as well as the temperature $T$ and the Hubble parameter $H$ remain approximately constant during the inflation stage. And when the Universe transits to the radiation domination stage, $I_{\mathrm{GW}}$ will be proportional to $e^{-N}$ since $T\propto e^{-N}$ and $H\propto e^{-2N}$. Therefore, $I_{\mathrm{GW}}(N)$ should sharply peaked at some e-fold, $N_{\mathrm{peak}}^{\mathrm{GW}}$, which satisfies $dI_{\rm GW}/dN=0$, or more explicitly:
$\left.\left[4+7\frac{d\ln{T(N)}}{dN}-\frac{d\ln{H(N)}}{dN}\right]\right|_{N=N_{\mathrm{peak}}^{\mathrm{GW}}}=0$ (3.9)
Thus, GW production is peaked around this maximum, and the integral in eq. (3.7) is dominated by contributions from the vicinity of $N_{\rm peak}^{\rm GW}$. This allows us to estimate the GW energy density at some late times for $N>N_{\rm peak}^{\rm GW}$
$\rho_{\rm GW}(N)\simeq e^{-4N}I_{\rm GW}(N_{\rm peak}^{\rm GW})\Delta N_{\mathrm{peak}}^{\rm GW}\quad(N>N_{\mathrm{peak}}^{\rm GW})$ (3.10)
@@
with $\Delta N_{\mathrm{peak}}^{\rm GW}\gtrsim 1$ being the half-width of $I_{\rm GW}(N)$.
The evolution of $I_{\rm GW}$ and $e^{4N}\rho_{\rm GW}$ are presented in figure 1. It can be clearly seen that most GW are produced around the peak of $I_{\rm GW}$, which occurs after the end of inflation but before the radiation-dominated epoch.
Figure 1: The comoving GW energy density $e^{4N}\rho_{\rm GW}$ and its production rate $I_{\rm GW}$ in WI with dissipation term $\Upsilon\propto T^{3}$ and $\Upsilon\propto T$, respectively. All curves are normalized to their respective maximum values. The blue line corresponds to $Q_{\mathrm{ini}}=0.01$ and the red line corresponds to $Q_{\mathrm{ini}}=1$. The black dashed line indicates the end of inflation ($\epsilon_{H}=1$), while the red and blue dashed lines correspond to the onset of radiation domination ($\epsilon_{H}=2$), respectively. Note that the reheating process is more efficient for $\Upsilon\propto T^{3}$ than for $\Upsilon\propto T$.
To better understand the freeze-in production of gravitons in WI, we compare it to the CI scenario. In CI, the production of freeze-in gravitons commences only after the Universe enters the radiation-dominated era. In contrast, gravitons can be produced over the entire inflationary epoch in WI, and as we have seen that most graviton energy is produced before the onset of radiation domination ($\epsilon_{H}=2$). Thus, we expect graviton production to be more efficient in WI. It is therefore natural to quantify the enhanced production in WI by comparing the total GW energy density in WI with that produced solely during the radiation-dominated epoch. Equivalently, this corresponds to comparing the gravitons production in WI with that in a CI scenario having the same reheating temperature $T_{\rm RH}$, which is defined as $T_{\rm RH}\equiv T(N_{\rm RH})$ with $N_{\rm RH}\equiv N(\epsilon_{H}=2)$.
It is then useful to normalize the GW energy density by defining the dimensionless quantity $\rho_{\rm GW}/s^{\frac{4}{3}}$ with $s=\frac{2\pi^{2}g_{*s}}{45}T^{3}$ the entropy density of the thermal bath [36]. Eq. (3.10) gives:
@@
$\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{e^{4(N_{\mathrm{peak}}^{\rm GW}-N)}}{M_{\rm pl}^{2}T^{4}(N)}\Delta N_{\mathrm{peak}}^{\rm GW}\times\frac{T^{7}(N_{\mathrm{peak}}^{\rm GW})}{H(N_{\mathrm{peak}}^{\rm GW})}\quad(N>N_{\mathrm{peak}}^{\rm GW})$ (3.11)
@@
For CI with the same reheating temperature $T_{\mathrm{RH}}$, by integrating eq. (3.7) from $N_{\mathrm{RH}}$ to infinity we can obtain
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI}}\simeq\left(\frac{2\pi^{2}g_{*s}}{45}\right)^{-\frac{4}{3}}\mathcal{C}_{1}\frac{T_{\rm RH}^{3}}{M_{\rm pl}^{2}H_{\rm RH}}$ (3.12)
where $H_{\rm RH}$ is the Hubble parameter at $N_{\rm RH}$.
@@
$\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]\bigg/\bigg[\frac{\rho_{\rm GW}}{s^{\frac{4}{3}}}\bigg]_{\mathrm{CI}}\simeq\frac{I_{\rm GW}(N_{\rm peak}^{\rm GW})}{I_{\rm GW}(N_{\rm RH})}\Delta N_{\mathrm{peak}}^{\rm GW}$ (3.13)
We conclude that GW are primarily generated around $N_{\rm peak}^{\rm GW}$, with their production enhanced in WI. Moreover, we expect a more pronounced enhancement for the WI model with $\Upsilon\propto T$ than for the $\Upsilon\propto T^{3}$ model, as the more efficient reheating process in the latter case leads to a significantly smaller value of the ratio $I_{\rm GW}(N^{\rm GW}_{\rm peak})\big/I_{\rm GW}(N_{\rm RH})$.